### v1

@@ 2.1 Scale Toward Exact Solutions with LAVA
Figure 2: Neural scaling laws enable breakthrough performance. a. LAVA delivers near-exact ground state energy for benzene as neural networks scale up. The neural scaling law here is described by $E-E_{\mathrm{expt}}\propto N_{p}^{-0.52}$, where $E$ denotes LAVA energies, $E_{\mathrm{expt}}$ denotes experiment-derived reference [20], $N_{p}$ is the number of parameters in the neural networks. We also include energies from NNVMC [9] and CCSD(T)/CBS [21] for comparison. b. Upper: Absolute energy errors of LAVA-SE, CCSDTQ5/CBS level W4-theory [22], CCSD(T)/CBS level W2-theory [22], and NNVMC. LAVA-SE is more accurate than 1 kJ/mol thresholds (purple shade), aligning closely with highly accurate W4 theory. Experimental references are based on atomization energies from ATcT (Active Thermochemical Tables) [22] and absolute atom energies from Chakravorty et al. [23]. Specifically, for atom F, we use our variationally lower result since it is more reliable. Details are in Supplementary Note 4.1. NNVMC uses LapNet [9] with the default model size. Lower: Neural scaling laws of LAVA on the same molecules, showing error reduction with model size in power-law trends. See Supplementary Note 2 for details. c. Cost scaling exponents for N2 as LAVA increases model size and CC methods increase excitation orders. For CC, data from Chan et al. [24] shows that CC’s computational effort scales as a power-law trend relative to $1/(E_{\mathrm{CC}}-E_{\mathrm{FCI}})$ under cc-pVDZ basis set. For LAVA, the total GPU time cost scales as a power-law relative to $1/(E_{\mathrm{LAVA}}-E_{\mathrm{SE}})$, where $E_{\mathrm{SE}}$ denotes the LAVA-SE result. LAVA maintains a nearly constant exponent along the $N_{2}$ dissociation curve, while CC’s performance deteriorates in the strongly correlated region. d. LAVA GPU hours as the number of electrons $N_{e}$ increases, with increasing model size for better accuracy. With model size fixed, runtime scales as $N_{e}^{2.1}$ (dotted lines). To ensure chemical accuracy, runtime scales as $N_{e}^{5.3}$ (the solid line).
@@ 2.2 Definitive Benchmarks beyond Experimental Limitations
Figure 3: Definitive LAVA benchmarks complement experiments. a. Upper: Reaction barriers of $\mathrm{C}_{4}\mathrm{H}_{4}$ derived from experiment and different theoretical methods. Bars represent the spread of reported values, corresponding to each line, within each method family. Stars indicate the improved experimental estimate and the best estimates from systematically improvable theories, namely CCSDTQ/aug-cc-pVTZ in CC, HCI/cc-pVTZ extrapolated to the FCI limit in selected CI, and LAVA in NNQMC. The dashed line shows the value predicted by LAVA, highlighting the consistency among those stars. Lower: Power-law scaling trends between the number of parameters $N_{p}$ and energy difference $E-E_{\mathrm{SE}}$, where $E_{\mathrm{SE}}$ are scaling-law extrapolation results. b. Upper main panel: The ground-state potential energy curve of N2 dissociation. The analytic potential functions, namely $\text{MLR}_{3}(9)$, $\text{MLR}_{4}(6,8)$, and $\text{EMO}_{2}(6)$, are fitted from vibrational levels $v=0-19$. Upper inset: PECs fitted from different analytic function forms and parameters show a large discrepancy at long bond lengths. Lower main panel: The energy difference between two ab initio calculations and three fitted analytic PECs. $\text{MLR}_{3}(9)$, the surrogate model of LAVA, is the new recommended benchmark. The $r_{12}$-MR-ACPF result is shifted together so that its results match high-fidelity experimental data in the equilibrium region (the shaded area).
@@
Next, we demonstrate LAVA on the potential energy curve of the nitrogen dimer, a prototypical and long-standing challenge due to its multireference character, and establish a new benchmark, MLR3(9), that surpasses the previous standard MLR4(6,8) in accuracy. The previous benchmark, labeled as MLR4(6,8) in Fig. 3.b, was fitted to experimental vibrational levels $v=0-19$ [39]. These vibrational levels provide a high-accuracy energy benchmark only in the near-equilibrium region, corresponding to bond lengths $r$ between 0.9 to 1.5 $\mathrm{\AA }$ (gray shaded region). There are few experimental results for vibrational levels $v$=20-25 [40], albeit with a high uncertainty. Measuring higher vibrational levels is challenging because the nitrogen dimer becomes unstable in the near-dissociation region.
@@
[…]
@@ 2.4 Beyond Energies
Figure 5: LAVA delivers reliable dipole moments. a. Dipole moments $\mu$ derived from our wavefunctions match experimental values $\mu_{\mathrm{expt}}$ [62]. For comparison, CCSD(T)/CBS results from Hait and Head-Gordon [63] are also visualized, which fall out of the experimental range for NO, O3, and F2O2. For all molecules, the SCAN density functional results [63] fail to match experimental references. b. Basis set convergence of dipole moment of F2O2 for HF and coupled-cluster methods, together with experimental benchmark and LAVA estimates for comparison. Dipole moments in Debyes are plotted against correlated-consistent basis sets (cc-pVXZ) and augmented counterparts (aug-cc-pVXZ) for $\mathrm{X}=\mathrm{D},\mathrm{T},\mathrm{Q},5$. CBS extrapolation scheme follows Supplementary Note 6.2. The missing points indicate calculations infeasible with available computational resources. Results of coupled-cluster methods exhibit strong basis set dependence and severe memory bottlenecks, while LAVA achieves experimental-level accuracy.
@@ 4.2 LAVA : the Lookahead Variational Algorithm
14:  $g\leftarrow-\frac{1}{B}\sum\limits_{i=1}^{B}\left\langle\mathrm{sgn}(E_{L}^{\prime}(x^{(i)}))\mathrm{sgn}(\nabla_{r^{(i)}}E_{L}^{\prime}(x^{(i)}))\right.,\left.\nabla_{\theta}\nabla_{r^{(i)}}\mathrm{ln}|\psi_{\theta}(x^{(i)})|\right\rangle\end{aligned}$
@@ 4.3 Neural Scaling Laws
where $\alpha$ and $\beta$ are system-dependent variables and $E_{0}$ is the exact ground state energy. We monotonically increase network width and the number of determinants during this scaling-up process. For 111111 systems with reliable experimental benchmarks, the average $r^{2}$ of ordinary least squares (OLS) regression on a logarithmic scale is larger than 0.950.950.95, and residual diagnostics (White test, $p>0.2$) reveals no significant heteroscedasticity. For the details, see Supplementary Note 2.
@@
where $\alpha_{v}$ and $\beta_{v}$ are system-dependent variables. Under the zero-variance principle of quantum Monte Carlo, our estimated energy approximates the exact ground state energy $E_{0}$ as $V\rightarrow 0$. Since variance estimation requires no exact reference data, unlike error measurement, verifying the relationship is possible for any system where LAVA calculations are available. We confirmed this relationship across various molecular systems by linear regression on a logarithmic scale, and the average $r^{2}$ is larger than 0.970.970.97. Still, OLS is used since the White test shows $p>0.2$. The detailed results are in Supplementary Note 2, and model architectures and training scheme are listed in Supplementary Note 1.1.
@@ 4.4 Extrapolation Scheme
Since theoretically $\mathop{\mathrm{lim}}\limits_{V\rightarrow 0}E=E_{0}$ and empirically $\mathop{\mathrm{lim}}\limits_{N_{p}\rightarrow+\infty}V=0$, we follow Fu et al. [80] and use $b$ as our asymptotic estimates. For various molecular systems, residual diagnostics (White test, $p>0.2$) support the use of ordinary least squares regression, and the average OLS $r^{2}$ for linear regression lines is greater than 0.990.990.99. For systems with experimental reference values, the errors in our extrapolation fall within the experimental uncertainty. Combining neural scaling laws and extrapolation enables the estimation of threshold computational resources to reach sub-kJ/mol accuracy.
@@ Acknowledgements
We thank Chenyang Li, Xuefei Xu, Yinan Shu, Donald G. Truhlar for the insight discussion. We thank Zigeng Huang, Qiming Sun, Xiaojie Wu, and the rest of the ByteDance Seed Group for their inspiring ideas and encouragement. We also thank Hang Li for his guidance and support. L.W. and D.H. are supported by National Science and Technology Major Project (2022ZD0114902) and National Science Foundation of China (NSFC92470123, NSFC62276005). W.A.G. thanks the US NSF (CBET 2311117) for support. J.C. is supported by the National Key R&D Program of China (2021YFA1400500) and National Science Foundation of China (12334003).

### v2

@@ 2.1 Scale Toward Exact Solutions with LAVA
Figure 2: Neural scaling laws enable breakthrough performance. a. LAVA delivers near-exact ground state energy for benzene as neural networks scale up. The neural scaling law here is described by $E-E_{\mathrm{expt}}\propto N_{p}^{-0.52}$, where $E$ denotes LAVA energies, $E_{\mathrm{expt}}$ denotes experiment-derived reference [20], $N_{p}$ is the number of parameters in the neural networks. We also include energies from NNVMC [9] and CCSD(T)/CBS [21] for comparison. b. Upper: Absolute energy errors of LAVA-SE, CCSDTQ5/CBS level W4-theory [22], CCSD(T)/CBS level W2-theory [22], and NNVMC. LAVA-SE is more accurate than 1 kJ/mol thresholds (purple shade), aligning closely with highly accurate W4 theory. Experimental references are based on atomization energies from ATcT (Active Thermochemical Tables) [22] and absolute atom energies from Chakravorty et al. [23]. Specifically, for atom F, we use our variationally lower result since it is more reliable. Details are in Supplementary Note 4.1. NNVMC uses LapNet [9] with the default model size. Lower: Neural scaling laws of LAVA on the same molecules, showing error reduction with model size in power-law trends. See Supplementary Note 2 for details. c. Cost scaling exponents for N2 as LAVA increases model size and $C$ methods increase excitation orders. For $C$, data from Chan et al. [24] shows that $C$’s computational effort scales as a power-law trend relative to $1/(E_{\mathrm{CC}}-E_{\mathrm{FCI}})$ under $c$-pVDZ basis set. For LAVA, the total GPU time cost scales as a power-law relative to $1/(E_{\mathrm{LAVA}}-E_{\mathrm{SE}})$, where $E_{\mathrm{SE}}$ denotes the LAVA-SE result. LAVA maintains a nearly constant exponent along the $N_{2}$ dissociation curve, while $C$’s performance deteriorates in the strongly correlated region. d. LAVA GPU hours as the number of electrons $N_{e}$ increases, with increasing model size for better accuracy. With model size fixed, runtime scales as $N_{e}^{2.1}$ (dotted lines). To ensure chemical accuracy, runtime scales as $N_{e}^{5.3}$ (the solid line).
@@ 2.2 Definitive Benchmarks beyond Experimental Limitations
Figure 3: Definitive LAVA benchmarks complement experiments. a. Upper: Reaction barriers of $\mathrm{C}_{4}\mathrm{H}_{4}$ derived from experiment and different theoretical methods. Bars represent the spread of reported values, corresponding to each line, within each method family. Stars indicate the improved experimental estimate and the best estimates from systematically improvable theories, namely CCSDTQ/aug-$c$-pVTZ in $C$, HCI/$c$-pVTZ extrapolated to the FCI limit in selected CI, and LAVA in NNQMC. The dashed line shows the value predicted by LAVA, highlighting the consistency among those stars. Lower: Power-law scaling trends between the number of parameters $N_{p}$ and energy difference $E-E_{\mathrm{SE}}$, where $E_{\mathrm{SE}}$ are scaling-law extrapolation results. b. Upper main panel: The ground-state potential energy curve of N2 dissociation. The analytic potential functions, namely $\text{MLR}_{3}(9)$, $\text{MLR}_{4}(6,8)$, and $\text{EMO}_{2}(6)$, are fitted from vibrational levels v=019v=0-19. Upper inset: PECs fitted from different analytic function forms and parameters show a large discrepancy at long bond lengths. Lower main panel: The energy difference between two ab initio calculations and three fitted analytic PECs. $\text{MLR}_{3}(9)$, the surrogate model of LAVA, is the new recommended benchmark. The $r_{12}$-MR-ACPF result is shifted together so that its results match high-fidelity experimental data in the equilibrium region (the shaded area).
@@
Next, we demonstrate LAVA on the potential energy curve of the nitrogen dimer, a prototypical and long-standing challenge due to its multireference character, and establish a new benchmark, MLR3(9), that surpasses the previous standard MLR4(6,8) in accuracy. The previous benchmark, labeled as MLR4(6,8) in Fig. 3.b, was fitted to experimental vibrational levels v=019v=0-19 [39]. These vibrational levels provide a high-accuracy energy benchmark only in the near-equilibrium region, corresponding to bond lengths $r$ between 0.9 to 1.5 $\mathrm{\AA }$ (gray shaded region). There are few experimental results for vibrational levels vv=20-25 [40], albeit with a high uncertainty. Measuring higher vibrational levels is challenging because the nitrogen dimer becomes unstable in the near-dissociation region.
@@
We now provide a definitive benchmark to resolve these discrepancies by performing LAVA calculations across various bond lengths. Furthermore, we fit an analytic potential curve in the Morse/long-range form [39], namely MLR3(9), based on LAVA data points which also reproduce the experimental vibrational levels vv=0-19. Details on this fitting procedure are given in the Supplementary Note 7. This new benchmark MLR3(9) retains the accuracy of the previous benchmark MLR4(6,8) [39] for the near-equilibrium region (0.9Å<r<1.5Å0.9\ $\mathrm{\AA }$<r<1.5\ $\mathrm{\AA }$) and the fully dissociation limit(r>4.0År>4.0\ $\mathrm{\AA }$), while improving the reliability around the near dissociation region. In addition, $r_{12}$-MR-ACPF [41, 45], an explicitly correlated multireference method with FCI/CBS accuracy, perfectly aligns with our new benchmark MLR3(9) within $1$ kJ/mol difference, after shifting down by about 5 mHa. However, $r_{12}$-MR-ACPF exhibits exponential scaling with respect to the size of reference space. This computational complexity necessitates careful optimization of both the reference space and basis set, thereby restricting its practical application to small molecular systems. Together, the reproduction of both the available experimental vibrational levels and the relative energies from $r_{12}$-MR-ACPF confirms the accuracy and reliability of our newly fitted benchmark MLR3(9) curve.
@@ 2.4 Beyond Energies
Figure 5: LAVA delivers reliable dipole moments. a. Dipole moments $\mu$ derived from our wavefunctions match experimental values $\mu_{\mathrm{expt}}$ [62]. For comparison, CCSD(T)/CBS results from Hait and Head-Gordon [63] are also visualized, which fall out of the experimental range for NO, O3, and F2O2. For all molecules, the SCAN density functional results [63] fail to match experimental references. b. Basis set convergence of dipole moment of F2O2 for HF and coupled-cluster methods, together with experimental benchmark and LAVA estimates for comparison. Dipole moments in Debyes are plotted against correlated-consistent basis sets ($c$-pVXZ) and augmented counterparts (aug-$c$-pVXZ) for $\mathrm{X}=\mathrm{D},\mathrm{T},\mathrm{Q},5$. CBS extrapolation scheme follows Supplementary Note 6.2. The missing points indicate calculations infeasible with available computational resources. Results of coupled-cluster methods exhibit strong basis set dependence and severe memory bottlenecks, while LAVA achieves experimental-level accuracy.
@@ 4.2 LAVA : the Lookahead Variational Algorithm
14:  $\begin{aligned} g\leftarrow-\frac{1}{B}\sum\limits_{i=1}^{B}\left\langle\mathrm{sgn}(E_{L}^{\prime}(x^{(i)}))\mathrm{sgn}(\nabla_{r^{(i)}}E_{L}^{\prime}(x^{(i)}))\right.,\left.\nabla_{\theta}\nabla_{r^{(i)}}\mathrm{ln}|\psi_{\theta}(x^{(i)})|\right\rangle\end{aligned}$
@@ 4.3 Neural Scaling Laws
where $\alpha$ and $\beta$ are system-dependent variables and $E_{0}$ is the exact ground state energy. We monotonically increase network width and the number of determinants during this scaling-up process. For $11$ systems with reliable experimental benchmarks, the average $r^{2}$ of ordinary least squares (OLS) regression on a logarithmic scale is larger than $0.95$, and residual diagnostics (White test, $p>0.2$) reveals no significant heteroscedasticity. For the details, see Supplementary Note 2.
@@
where $\alpha_{v}$ and $\beta_{v}$ are system-dependent variables. Under the zero-variance principle of quantum Monte Carlo, our estimated energy approximates the exact ground state energy $E_{0}$ as $V\rightarrow 0$. Since variance estimation requires no exact reference data, unlike error measurement, verifying the relationship is possible for any system where LAVA calculations are available. We confirmed this relationship across various molecular systems by linear regression on a logarithmic scale, and the average $r^{2}$ is larger than $0.97$. Still, OLS is used since the White test shows $p>0.2$. The detailed results are in Supplementary Note 2, and model architectures and training scheme are listed in Supplementary Note 1.1.
@@ 4.4 Extrapolation Scheme
Since theoretically $\mathop{\mathrm{lim}}\limits_{V\rightarrow 0}E=E_{0}$ and empirically $\mathop{\mathrm{lim}}\limits_{N_{p}\rightarrow+\infty}V=0$, we follow Fu et al. [80] and use $b$ as our asymptotic estimates. For various molecular systems, residual diagnostics (White test, $p>0.2$) support the use of ordinary least squares regression, and the average OLS $r^{2}$ for linear regression lines is greater than $0.99$. For systems with experimental reference values, the errors in our extrapolation fall within the experimental uncertainty. Combining neural scaling laws and extrapolation enables the estimation of threshold computational resources to reach sub-kJ/mol accuracy.
@@ Acknowledgements
We thank Chenyang Li, Xuefei Xu, Yinan Shu, Donald G. Truhlar for the insight discussion. We thank Zigeng Huang, Qiming Sun, Xiaojie Wu, and the rest of the ByteDance Seed Group for their inspiring ideas and encouragement. We also thank Hang Li for his guidance and support. L.W. is supported by National Science and Technology Major Project (2022ZD0114902) and National Science Foundation of China (NSFC92470123, NSFC62276005). D.H. is supported by National Science Foundation of China (NSFC62376007). W.A.G. thanks the US NSF (CBET 2311117) for support. J.C. is supported by the National Key R&D Program of China (2021YFA1400500) and National Science Foundation of China (12334003).